OpenStudy (anonymous):

Can someone please solve this problem.Find K so that the line through (k-1,k+2) and (4,-1) is perpendicular to the line through (-3,2) and (2,5)

OpenStudy (campbell_st):

ok... what did you get as the slope of the passing through (-3, 2) and (2, 5)..?

OpenStudy (anonymous):

I still dont understand it

OpenStudy (campbell_st):

so the slope of the line through (-3, 2) and (2, 5) is (5 -2)/(2 - -3) = 1 so then the perpendicular is the negative reciprocal to the slope of the line through (k -1, k + 2) and (4, -1) is m = -1 then using the formula for the slope of a line \[-1 = \frac{k + 2 - -1}{k -1 - 5} \] which can be simplified to \[-1 = \frac{k + 3}{k - 6}\] all you need to do is solve for k